Use the arithmetic sequence of numbers 2, 4, 6, 8, 10, what is the sum of the first 30 terms? I came up with 930, am I correct?
Also how do I type in the term or subscript? Ex: a1 a2 a3.....

Answer by AnlytcPhil(1806)


Use the arithmetic sequence of numbers

2, 4, 6, 8, 10,

what is the sum of the first 30 terms?
I came up with 930, am I correct? 

2nd term - 1st term =  4 - 2 = 2
3rd term - 2nd term =  6 - 4 = 2
4th term - 3rd term =  8 - 6 = 2
5th term - 4th term = 10 - 8 = 2

These are all 2, so the common difference, d = 2.

The sum of the first n terms is given by the formula

Sn = n[2a1 + (n-1)d]/2

n = 30, a1 = 2, d = 2

S30 = 30[2(2) + (30-1)(2)]/2

S30 = 30[4 + (29)(2)]/2

S30 = 30(4 + 58)/2

S30 = 30(62)/2 

S30 = 1860/2

S30 = 930

Looks like you're right!

-------------------------------------------

Also how do I type in the term or subscript? 
Ex: a1 a2 a3.....

With HTML tags.  To type

a1 

you have to type 

a< sub>1< /sub>

except you close those spaces after the < 's. I didn't 
here because if I had you'd have only seen a1.
I don't type the tags individually.  That takes too long.
I put the tags in after I finish, using 'cut and paste'.

Edwin